# mystylerightline=true, innerleftmargin=10, innerrightmargin=10, outerlinewidth=3pt, topline=false, rightline=false, bottomline=false, skipabove=skipbelow=false

### v1

@@ 9      $ppl\leftarrow\texttt{compute\_ppl}(\Theta^{{}^{\prime}}(S_{ppl}))$
10      =$\mathcal{B}=\mathcal{B}\cup\{(s,bias)\}$
@@ 11      $\mathcal{P}=\mathcal{P}\cup\{(s,ppl)\}$
Input: Unfair LLM $\Theta$, DNNs to predict bias and perplexity $[\theta_{bias},\theta_{ppl}]$, min. and max. number of attention heads to drop $[n_{l},n_{u}]$, and Timeout $time\_limit$
@@
10      $T\leftarrow$ update_temperature($T_{0}$, $curr\_time()$)
@@ 11      $s_{i}\leftarrow$ generate_state($s,n_{l},n_{u}$)
12      $cost_{i}$leftarrow $\theta_{bias}(s_{i})+\theta_{ppl}(s_{i})$
@@ 5. Experiments
The architecture for the surrogate DNNs is shown in Table 2. We find that small DNNs with two hidden layers can effectively predict a state’s bias and perplexity in the search space. The size of the input layer is the number of attention heads in the LLM. The output layer has one neuron. For the smaller LLMs ($<1\text{B}$ parameters) with fewer attention heads, we find that hidden layers with more neurons tend to overfit. Thus, we opt to use small hidden layers with 64 and 32 neurons. For the bigger LLMs, hidden layers with 256 and 128 neurons perform better. The size of the datasets used to train the surrogate DNNs is highlighted in Table 2. We use a train-validation split of 95%-5%. The DNNs are trained till the validation split error starts to increase. During the dataset creation in Algorithm 1, we set the upper bound on the number of attention heads dropped $n_{u}$ to up to $10\%-20\%$ of the total attention heads in the LLM. We set $\eta_{bias}\in[0.1,0.2]$ and $\eta_{ppl}=1$. Higher values of $\eta_{bias}$ are chosen for the smaller LLMs since their inference times are faster. We use the entire WikiText-2 validation split as it only contains $2,461$ text sequences. Additionally, dropping attention heads increases perplexity and reduces bias since meaningless and random text has a low bias by default. Therefore, we need the perplexity DNN $\theta_{ppl}$ to be as accurate as possible to ensure that the SA search returns states with low bias and perplexity. Therefore, to ensure high data quality for the perplexity samples, we used the entire WikiText-2 validation split. On the other hand, the HolisticBias validation dataset has $10,268$ gender bias prompts, and we only use a fraction of them to speed up the data collection process. We perform experiments on a shared GPU cluster with variable GPU availability depending on server load. We note that the dataset collection process is offline. Instead of setting a time limit for Algorithm 1, we collect at least 25,000 samples to train the DNNs. Once the surrogate DNNs are trained, they will be used in the online phase for the SA search.
[…]
All algorithms were implemented in python=3.10. The surrogate DNNs were implemented with torch==2.2.2. The LLMs were implemented using transformers==4.29.0 and torch==2.4.1.
@@
Data pre-processing. Regression with DNNs is effective if the predicted values lie in $[0,1]$ as machine learning models operate well on small floating point numbers. However, the bias and perplexity of LLMs may take values greater than 1. We observe that the maximum bias in the dataset $\mathcal{B}$ we create to train the bias DNN $\theta_{bias}$ is between $0.9-1.1$ depending on the LLM. Since the maximum is already close to 1, we simply scale the bias by dividing all the values by the maximum. The perplexity, on the other hand, is much harder to pre-process. Table 3 highlights the statistics of the perplexity in our dataset. We observe that perplexity takes large floating point numbers and has high variance. The dataset is quite noisy, and during our initial experiments, we observed that simply scaling the perplexity by the maximum leads to sub-optimal results. For example, if we were to scale the perplexity of the GPT-Neo-125M dataset, the baseline perplexity would be $\approx 2.9\times 10^{-11}$. Such minuscule values make it difficult to differentiate between states that have low perplexities. For example, states with perplexity 37 and 45 (which is $\approx 3.5\times 10^{-11}$ after scaling) would be treated similarly by the DNN. Additionally, it is difficult to train DNNs to predict such small values reliably. One option to reduce variance would be to remove all data points with high perplexity from the dataset. However, this is undesirable as we want our SA search to avoid such states that disproportionately degrade model utility. Our solution, therefore, is to clamp the perplexity such that the standard deviation is less than a pre-defined threshold. Formally,
@@
Biased Subsampling. In Algorithm 1, we subsample a fraction $\eta_{bias}$ of the gender bias prompts from the HolisticBias validation dataset to prompt the LLM at each iteration. The gender bias group has multiple subgroups denoted by set $G$. For example, some of the members of $G$ are binary, queer, transgender, etc. The number of prompts for each subgroup of gender bias in the HolisticBias dataset is not uniform. For example, among all the gender bias prompts, 40.0% of them belong to the non-binary subgroup while 2.7% belong to the binary sub-group. Such imbalances are a common occurrence in datasets used for fairness evaluations. While training $\theta_{bias}$, we observe that datasets generated using random sampling that does not account for the underlying distribution lead to poor results. The prediction of $\theta_{bias}$ trained on randomly sampled subsets does not reflect the true bias of the LLM on the entire HolisticBias validation dataset. Therefore, we perform biased subsampling to sample subsets. We ensure that the subgroups in each subset are similar to the distribution in the entire dataset, i.e., 40% of the prompts in the subsampled dataset belong to non-binary and 2.7% belong to binary. Similarly, the distribution is followed for the other subgroups (transgender, queer, etc.).
@@ 4. RQ4
    How does reducing gender bias with Attention Pruning affect other social biases?
5.0.1. RQ1: Effectiveness of surrogate DNNs at estimating bias and perplexity characteristics of attention heads
@@
5.0.2. RQ2: Effectiveness of Attention Pruning compared to state-of-the-art pruning strategies
Table 6. Bias and perplexity of LLMs before and after bias mitigation techniques are applied. Our tool, AP, outperforms the state-of-the-art attention head pruning strategies.
@@ 5.0.2. RQ2: Effectiveness of Attention Pruning compared to state-of-the-art pruning strategies
We rigorously evaluate our approach against four pruning strategies. The FASP algorithm has two hyperparameters: $\alpha$ and $\gamma$, where $\alpha$ denotes the ratio of heads that are pruned and $\gamma$ denotes the ratio of heads important for utility that are not pruned. The authors perform an exhaustive search for $\gamma$ and determine the optimal value is $0.6$ for GPT-Neo-1.3B and $0.3$ for the other LLMs. However, for $\alpha$, they do not provide a quantitative analysis. They vary $\alpha\in[0.02,0.2]$ in increments of 0.02 and show how the bias and perplexity change compared to general-purpose pruning strategies ($\alpha$ for these strategies, too, denotes the ratio of heads that are pruned). For an objective comparison with other approaches, we perform an exhaustive search for $\alpha\in[0.02,0.2]$ in increments of 0.02 and pick the value with the lowest bias. AP has three important hyperparameters: $\epsilon,n_{l},n_{u}$. We set $n_{l}=2$ except for GPT-Neo-125M, where $n_{l}=n_{u}$; i.e., we exactly prune a fixed number of heads during the search for GPT-Neo-125M. We observe that this strategy yields better results for GPT-Neo-125M. For other models, we set $n_{u}$ to be equal to a fixed ratio $\eta$ of the total number of heads i.e., $n_{u}=\eta\times N$, where $N$ is the total number of heads in the LLM. Once we have identified a good set of hyperparameters for FASP and AP, we generate text with pruned LLMs using three different seeds to sample tokens and report the mean and standard deviation of the bias. We set the time limit for the SA search to 3 hours. The perplexity is independent of the seed since it is always evaluated on the WikiText-2 test dataset. For the random pruning strategy, we randomly prune up to $20\%$ of the heads and report the best value observed.
Table 6 presents the bias and perplexity of the baseline LLM, AP, and the other strategies. The values in bold correspond to the best results. We see that in 4/6 cases, AP finds states with better bias and perplexity than FASP. The largest absolute and relative improvement in the bias of $0.18$ and $40.8\%$ is observed for GPT-J-6B. For GPT-Neo-125M, AP’s bias is worse by $\approx 0.02$, but the perplexity is better by $\approx 6$. For Llama-2-7B, AP’s bias is better by $\approx 0.03$ and the perplexity is worse by $\approx 0.3$. Overall, in most cases, AP can reduce bias more while ensuring better perplexity. The other strategies, which do not account for bias, are worse than both AP and FASP222 Gradient Magnitude for Llama-2-7B is left blank as its implementation in Huggingface does not provide an option for computing the gradient of masked heads..
AP can find states with better bias and perplexity since it explores regions of the search space that FASP cannot. FASP cannot explore certain parts of the search space since it does not account for the non-linear relationship among attention heads.
@@
5.0.3. RQ3: Design Considerations of Attention Pruning in mitigating unfairness in LLMs
@@
  • The cost function in SA is a linear weighted combination of the two surrogate DNNs. The hyperparameter $\epsilon$ determines how much weight to give to each DNN. Higher values of $\epsilon$ assign more weight to the bias DNN, prioritizing the reduction of bias over the loss of perplexity during the SA search. To visualize the bias-perplexity tradeoff in LLMs, we vary $\epsilon\in[0.3,0.7]$ during the SA search for four LLMs. Figure 2 shows plots of bias versus perplexity of the states for different epsilon values. States with low bias and high perplexity occur in the upper left region of the graph. States with high bias and low perplexity occur in the lower right region. From the plots, with the exception of outliers that occur due to the randomness of SA, we see that the smaller bubbles that represent lower values of $\epsilon$ occur in the lower right region. Similarly, the larger bubbles moving towards the upper right region represent larger values of $\epsilon$.
  • During the creation of the dataset for training the surrogate DNNs, increasing $\eta_{bias}$ and $\eta_{ppl}$ will use larger portions of the HolisticBias and WikiText-2 datasets while collecting the bias and perplexity statistics. This can lead to more accurate DNN predictions. Similarly, using more data samples could lead to better DNN predictions.
  • Decreasing $n_{l}$ and increasing $n_{u}$ during the SA search will increase the search space size. Increasing $n_{u}$ is equivalent to increasing $\eta$, which is the maximum ratio of attention heads that can be pruned; i.e., $n_{u}=\eta\times N$. A larger search space can benefit the SA search, provided that the total running time increases since it has more states to explore.
  • Increasing the running time ($time\_limit$) of the SA search may lead to better states as the search can explore more states.
@@
Answer RQ3: Using larger portions of the bias and perplexity datasets and collecting more data points to train the surrogate DNNs can increase their accuracy. Increasing the size of the search space and running SA longer may lead to better solutions. Increasing the weightage given to the bias DNN in the cost function reduces bias at the expense of perplexity.
5.0.4. RQ4: Effect of reducing gender bias on other social biases
@@
One of the possible unintended side effects of mitigating one form of bias is that it may exacerbate other forms of bias in the LLM. The HolisticBias metric contains prompts for 13 social biases, and we primarily focus on mitigating gender bias. Similar to Zayed et al. (2024), to understand how the other social biases are affected by AP, we pick 4 other social biases and evaluate them after repairing gender bias. Specifically, we pick the best subset of attention heads AP found for gender bias, prune the LLMs and finally evaluate the repaired LLM on the other biases. Table 7 highlights the effect on other biases before and after repairing gender bias. Note that the perplexity remains identical to the values reported in Table 6 as the perplexity is always evaluated with the WikiText-2 test dataset. We observe that, across the board, all other social biases decrease for all LLMs. We observe consistent improvement in all cases, even for biases like age, which are low by default. The highest absolute improvement in bias of 0.29 is observed for GPT-Neo-125M for sexual orientation bias. The highest relative improvement of $65.21\%$ is observed for GPT-J-6B in the age bias category.
Answer RQ4: Reducing gender bias with Attention Pruning positively affects other social biases. We observe that other social biases in all LLMs are reduced by up to $65\%$.

### v2

@@ 9 $ppl\leftarrow\texttt{compute\_ppl}(\Theta^{{}^{\prime}}(S_{ppl}))$
10 $\mathcal{B}=\mathcal{B}\cup\{(s,bias)\}$
@@ 11 $\mathcal{P}=\mathcal{P}\cup\{(s,ppl)\}$
Input: Unfair LLM $\Theta$, DNNs to predict bias and perplexity $\theta_{bias},\theta_{ppl}]$, min. and max. number of attention heads to drop $n_{l},n_{u}]$, and Timeout time_limittime\_limit
@@
10 $T\leftarrow$ update_temperature($T_{0}$, curr_time()curr\_time())
@@ 11 $s_{i}\leftarrow$ generate_state($s,n_{l},n_{u}$)
12 $cost_{i}$ $\leftarrow$ $\theta_{bias}(s_{i})+\theta_{ppl}(s_{i})$
@@ 5. Experiments
The architecture for the surrogate DNNs is shown in Table 2. We find that small DNNs with two hidden layers can effectively predict a state’s bias and perplexity in the search space. The size of the input layer is the number of attention heads in the LLM. The output layer has one neuron. For the smaller LLMs ($1\text{B}$ parameters) with fewer attention heads, we find that hidden layers with more neurons tend to overfit. Thus, we opt to use small hidden layers with 64 and 32 neurons. For the bigger LLMs, hidden layers with 256 and 128 neurons perform better. The size of the datasets used to train the surrogate DNNs is highlighted in Table 2. We use a train-validation split of 95%-5%. The DNNs are trained till the validation split error starts to increase. During the dataset creation in Algorithm 1, we set the upper bound on the number of attention heads dropped $n_{u}$ to up to $10\%-20\%$ of the total attention heads in the LLM. We set $\eta_{bias}\in[0.1,0.2]$ and $\eta_{ppl}=1$. Higher values of $\eta_{bias}$ are chosen for the smaller LLMs since their inference times are faster. We use the entire WikiText-2 validation split as it only contains $2,461$ text sequences. Additionally, dropping attention heads increases perplexity and reduces bias since meaningless and random text has a low bias by default. Therefore, we need the perplexity DNN $\theta_{ppl}$ to be as accurate as possible to ensure that the SA search returns states with low bias and perplexity. To ensure high data quality for the perplexity samples, we used the entire WikiText-2 validation split. On the other hand, the HolisticBias validation dataset has $10,268$ gender bias prompts, and we only use a fraction of them to speed up the data collection process. We perform experiments on a shared GPU cluster with variable GPU availability depending on server load. We note that the dataset collection process is offline. Instead of setting a time limit for Algorithm 1, we collect at least 25,000 samples to train the DNNs. Once the surrogate DNNs are trained, they will be used in the online phase for the SA search.
Technical Details. The dataset collection to train the surrogate DNNs and inference on the LLMs was performed on a shared GPU cluster running Ubuntu 20.04.6 LTS with an Intel(R) Xeon(R) Gold 6336Y CPU @ 2.40GHz, 10 NVIDIA RTX A6000 GPUs, and 512GB RAM. The SA runs were performed on a desktop running Ubuntu 22.04.3 LTS with an Intel(R) Core(TM) i7-7700 CPU @ $3.60\text{\,}\mathrm{G}\mathrm{H}\mathrm{z}$ processor, $32\text{\,}\mathrm{G}\mathrm{B}$ RAM, and a $1\text{\,}\mathrm{T}\mathrm{B}$ HDD.
All algorithms were implemented in python=3.10. The surrogate DNNs were implemented with torch==2.2.2. The LLMs were implemented using transformers==4.29.0 and torch==2.4.1. We open-source our implementations [^2] https://bitbucket.org/psu_soslab/attention_pruning.
@@
Data pre-processing. Regression with DNNs is effective if the predicted values lie in $[0,1]$ as machine learning models operate well on small floating point numbers. However, the bias and perplexity of LLMs may take values greater than 1. We observe that the maximum bias in the dataset $\mathcal{B}$ we create to train the bias DNN $\theta_{bias}$ is between 0.91.10.9-1.1 depending on the LLM. Since the maximum is already close to 1, we simply scale the bias by dividing all the values by the maximum. The perplexity, on the other hand, is much harder to pre-process. Table 3 highlights the statistics of the perplexity in our dataset. We observe that perplexity takes large floating point numbers and has high variance. The dataset is quite noisy, and during our initial experiments, we observed that simply scaling the perplexity by the maximum leads to sub-optimal results. For example, if we were to scale the perplexity of the GPT-Neo-125M dataset, the baseline perplexity would be $\approx 2.9\times 10^{-11}$. Such minuscule values make it difficult to differentiate between states that have low perplexities. For example, states with perplexity 37 and 45 (which is $\approx 3.5\times 10^{-11}$ after scaling) would be treated similarly by the DNN. Additionally, it is difficult to train DNNs to predict such small values reliably. One option to reduce variance would be to remove all data points with high perplexity from the dataset. However, this is undesirable as we want our SA search to avoid such states that disproportionately degrade model utility. Our solution, therefore, is to clamp the perplexity such that the standard deviation is less than a pre-defined threshold. Formally,
@@
Biased Subsampling. In Algorithm 1, we subsample a fraction $\eta_{bias}$ of the gender bias prompts from the HolisticBias validation dataset to prompt the LLM at each iteration. The gender bias group has multiple subgroups denoted by set $G$. For example, some of the members of $G$ are binary, queer, transgender, etc. The number of prompts for each subgroup of gender bias in the HolisticBias dataset is not uniform. For example, among all the gender bias prompts, 40.0% of them belong to the non-binary subgroup while 2.7% belong to the binary sub-group. Such imbalances are a common occurrence in datasets used for fairness evaluations. Because only 2.7% of the population is non-binary, a simple random sample—unless it is quite large—has a substantial chance of including few or even no non-binary individuals. For example, a purely random draw of 30 samples contains zero non-binary participants nearly half the time. Random sampling, therefore, tends to under-represent this subgroup. While training $\theta_{bias}$, we observe that datasets generated using random sampling that does not account for the underlying distribution lead to poor results. The prediction of $\theta_{bias}$ trained on randomly sampled subsets does not reflect the true bias of the LLM on the entire HolisticBias validation dataset. Therefore, we perform biased subsampling to sample subsets. We ensure that the subgroups in each subset are similar to the distribution in the entire dataset, i.e., 40% of the prompts in the subsampled dataset belong to non-binary and 2.7% belong to binary. Similarly, the distribution is followed for the other subgroups (transgender, queer, etc.).
@@ 4. RQ4
    How does reducing gender bias with Attention Pruning affect other social biases and can Attention Pruning generalize to other biases?
RQ1: Effectiveness of surrogate DNNs at estimating bias and perplexity characteristics of attention heads
@@
RQ2: Effectiveness of Attention Pruning compared to state-of-the-art pruning strategies
Table 6. Gender Bias and perplexity of LLMs before and after bias mitigation techniques are applied. Our tool, AP, outperforms the state-of-the-art attention head pruning strategies. The text in bold highlights the best bias and perplexity of fairness-aware pruning strategies that minimize bias while minimizing perplexity degradation.
@@
We evaluate our approach against four pruning strategies. The FASP algorithm has two hyperparameters: $\alpha$ and $\gamma$, where $\alpha$ denotes the ratio of heads that are pruned and $\gamma$ denotes the ratio of heads important for utility that are not pruned. The authors performed an exhaustive search for $\gamma$ and determined that the optimal value is $0.6$ for GPT-Neo-1.3B and $0.3$ for the other LLMs. However, for $\alpha$, they do not provide a quantitative analysis. They vary $\alpha\in[0.02,0.2]$ in increments of 0.02 and show how the bias and perplexity change compared to general-purpose pruning strategies ($\alpha$ for these strategies, too, denotes the ratio of heads that are pruned). For an objective comparison with other approaches, we perform an exhaustive search for $\alpha\in[0.02,0.2]$ in increments of 0.02 and pick the value with the lowest bias. AP has three important hyperparameters: $\epsilon,n_{l},n_{u}$. We set $n_{l}=2$ except for GPT-Neo-125M, where $n_{l}=n_{u}$; i.e., we exactly prune a fixed number of heads during the search for GPT-Neo-125M. We observe that this strategy yields better results for GPT-Neo-125M. For other models, we set $n_{u}$ to be equal to a fixed ratio $\eta$ of the total number of heads i.e., $n_{u}=\eta\times N$, where $N$ is the total number of heads in the LLM. Once we have identified a good set of hyperparameters for FASP and AP, we generate text with pruned LLMs using three different seeds to sample tokens and report the mean and standard deviation of the bias. We set the time limit for the SA search to 3 hours. The perplexity is seed-independent since it is always evaluated on the WikiText-2 test dataset. For the random pruning strategy, we randomly prune up to $20\%$ of the heads and report the best value observed.
Table 6 presents the bias and perplexity of the baseline LLM, AP, and the other strategies. The values in bold correspond to the best results. We see that in 4/6 cases, AP finds states with better bias and perplexity than FASP. The largest absolute and relative improvement in the bias of $0.18$ and $40.8\%$ is observed for GPT-J-6B. For GPT-Neo-125M, AP’s bias is worse by $\approx 0.02$, but the perplexity is better by $\approx 6$. For Llama-2-7B, AP’s bias is better by $\approx 0.03$ and the perplexity is worse by $\approx 0.3$. Overall, in most cases, AP can reduce bias more while ensuring better perplexity. The other strategies, which do not account for bias, are worse than both AP and FASP [^3] Gradient Magnitude for Llama-2-7B is left blank as its implementation in Huggingface does not provide an option for computing the gradient of masked heads..
@@
RQ3: Design Considerations of Attention Pruning in mitigating unfairness in LLMs
@@
  • The cost function in SA is a linear weighted combination of the two surrogate DNNs. The hyperparameter $\epsilon$ determines how much weight to give to each DNN. Higher values of $\epsilon$ assign more weight to the bias DNN, prioritizing the reduction of bias over the loss of perplexity during the SA search. To visualize the bias-perplexity tradeoff in LLMs, we vary $\epsilon\in[0.3,0.7]$ during the SA search for four LLMs. Figure 2 shows plots of bias versus perplexity of the states for different epsilon values. States with low bias and high perplexity occur in the upper left region of the graph. States with high bias and low perplexity occur in the lower right region. From the plots, except for outliers that occur due to the randomness of SA, we see that the smaller bubbles that represent lower values of $\epsilon$ occur in the lower right region. Similarly, the larger bubbles moving towards the upper right region represent larger values of $\epsilon$.
@@
  • During the creation of the dataset for training the surrogate DNNs, increasing $\eta_{bias}$ and $\eta_{ppl}$ will use larger portions of the HolisticBias and WikiText-2 datasets. This can lead to more accurate DNN predictions as larger subsets tend to accurately represent the true bias/perplexity of the entire dataset. To analyze how accurately the subsets represent the true bias, we sample 25 subsets each for different values of $\eta_{bias}\in[0.1,0.9]$ and plot the error in bias. Figure 3 shows a plot of the error versus $\eta_{bias}$. The dotted line is a regression line fitting all points. Overall, we observe a positive trend where the error decreases as $\eta_{bias}$ is increased i.e., larger portions of the validation dataset are used to compute the bias of each sampled subset.
    Figure 3. Absolute error of bias of subsets sampled with varying ratios $\eta_{bias}$ compared to true bias of entire validation dataset with GPT-J-6B.
  • Increasing the running time (time_limittime\_limit) of the SA search may lead to better states as the search can explore more states. To analyze the effect of running time on the cost and bias/perplexity of best states found, we run the SA search for 5.5 hours with $\epsilon=0.5$ and record the best result at 30-minute intervals. Figure 4 highlights the results for GPT-J-6B. From Figure 4(a), we see that the cost of the best state is a decreasing function. The cost is guaranteed to remain the same or decrease over time as the SA search explores more regions of the search space. However, bias/perplexity has a more complicated relationship. From Figure 4(b), we see that bias and perplexity alternate over time, eventually ending with the bias improving more at the expense of perplexity. We observe this because the cost function is a weighted linear function of bias and perplexity. From Figure 2, we see that the best option to control the bias/perplexity trade-off is to vary $\epsilon$.
    (a) Cost of best state
    (b) Bias/Perplexity of best state
    Figure 4. Effect of running time on the SA cost and bias/perplexity for GPT-J-6B
Answer RQ3: Using larger portions of the bias and perplexity datasets leads to more accurate estimates of bias/perplexity that can improve the DNN performance. Increasing the weightage given to the bias DNN in the cost function reduces bias at the expense of perplexity. Running SA longer may lead to better solutions.
RQ4: Effect of reducing gender bias on other social biases and generalizability
One of the possible unintended side effects of mitigating one form of bias is that it may exacerbate other forms of bias in the LLM. The HolisticBias metric contains prompts for 13 social biases, and we primarily focus on mitigating gender bias. Similar to Zayed et al. (2024), to understand how the other social biases are affected by AP, we pick 4 other social biases and evaluate them after repairing gender bias. We pick the best subset of attention heads AP found for gender bias, prune the LLMs, and finally evaluate the repaired LLM on the other biases. Table 7 highlights the effect on other biases before and after repairing gender bias. Note that the perplexity remains identical to the values reported in Table 6 as the perplexity is always evaluated with the WikiText-2 test dataset. We observe that all other social biases decrease for all LLMs. We see a consistent improvement in all cases, even for biases such as age, which are low by default. The highest absolute improvement in bias of 0.29 is observed for GPT-Neo-125M for sexual orientation bias. The highest relative improvement of $65.21\%$ is observed for GPT-J-6B in the age bias category.
@@
To ensure that AP generalizes beyond gender bias, we experiment with directly minimizing racial bias. This is distinct from Table 7, where we determine if reducing gender bias has a positive impact on other biases. Here, our goal is to minimize racial bias using AP directly. From Table 8, we see that AP effectively reduces racial bias in LLMs. Overall, comparing the values in Table 7 and Table 8, we see that directly reducing racial bias yields better results than reducing racial bias via gender bias.
Table 8. Racial Bias and Perplexity of LLMs before and after Attention Pruning is applied.
Model Baseline AP
Bias PPL Bias PPL
Distilgpt-2 0.529 62.325 0.356 73.157
GPT-2 0.518 43.588 0.336 58.461
GPT-Neo-125M 0.448 35.628 0.333 40.477
GPT-Neo-1.3B 0.463 17.422 0.375 18.427
GPT-J-6B 0.496 11.695 0.342 13.461
Llama-2-7B 0.458 6.781 0.363 7.375
Answer RQ4: Reducing gender bias with Attention Pruning positively affects other social biases. We observe that other social biases in all LLMs are reduced by up to $65\%$. Furthermore, AP is also effective at directly reducing other social biases.
Table 9. Qualitative example of bias in GPT-2 among the sex and cisgender subgroups. AP promotes fairer responses by ensuring they have similar toxicity across different subgroups. Sentences are truncated to 40 tokens.
Prompt Method Text Generation Toxicity
“I don’t like intersex boys. Baseline I’m not going to be a mother to any of them. I don’t support them. I’m open to having kids with girls that I know that are different from their biological 0.402
FASP You know, how many times can he say?” She says, “We already know that there are women who can be happy on this spectrum! Even if just a quarter of that spectrum is just fine 0.036
AP I’m not going to hold up my door.” As one woman told the BBC News channel, “I’ve been told that some of his other clients are now coming here, and that these” 0.097
“I don’t like cisgender boys. Baseline It is only by doing something that they become more and more capable of making decisions,” says Taylor, who was in her mid-50s when they went to the school when it first began. 0.025
FASP Let’s add that he’s out there and I’m happy that they’re taking on him in public life. Is he a threat? Sure, he’s very much a threat at this point.” 0.16
AP It is only half true because I believe in a system of privilege.” “To have a body like that, what he believes in and what he wants, that ‘This is my body,’ 0.03